# https://www.mmlab-ntu.com/project/edgesam

### v2

@@ 1 Introduction
In this study, we explore the feasibility of deploying the Segment Anything Model (SAM) [24] directly on edge devices, such as smartphones, to enable real-time interactive segmentation and facilitate its integration in various downstream tasks. Addressing this challenge is non-trivial and non-achievable with the usual engineering efforts of compressing the model. First, SAM was not designed for on-device deployment. It has a total parameter count of 641M and requires 2735 GFLOPs under its fixed $1024\times 1024$ input resolution. We failed to run SAM on an iPhone 14 due to its large computation and memory consumption. Even on an NVIDIA 2080 Ti, the throughput is just four images per second.
Figure 1: Figure (a) shows the overview of EdgeSAM. We first apply encoder-only KD between the output feature of the image encoder of SAM and EdgeSAM. Then we adopt prompt-in-the-loop KD, which interactively samples point prompts from the wrongly segmented areas. Note that, the initial prompt can also be a box. In (b), we show its throughput and the mIoU performance on the COCO and LVIS datasets with box prompts compared with SAM, MobileSAM, and EfficientSAM.
@@
A seemingly straightforward solution to speed up is to replace SAM’s huge ViT-based image encoder with a more compact version, as demonstrated by MobileSAM [61]. This approach indeed leads to a considerable speed boost, approximately 23 times faster, but it also substantially reduces the performance. For instance, the mask mIoU on the COCO dataset [31] drops from 77.3 to 74.4 with ground-truth boxes as the box prompts. Moreover, when deployed on edge devices, the speed of MobileSAM is still far from being real-time. For example, on iPhone 14, its throughput is only five images per second. A more recent work, EfficientSAM [57] follows the same architecture as MobileSAM but further leverages the masked image pre-training to boost the performance. However, it runs at the same speed as MobileSAM and consumes huge computational costs during training.
@@
In our extensive empirical investigations, it became unexpectedly clear that the key factor in knowledge distillation from SAM is not the incorporation of losses specifically tailored for dense tasks [32, 44] or query-based detectors [21, 7, 5]. Rather, it is the strategic selection of prompts during the distillation process that holds paramount importance. Addressing this, we introduce a novel approach: a dynamic prompt-in-the-loop strategy. This technique effectively enables the student model to assimilate the intricate knowledge encapsulated within SAM. It involves actively aligning the student model with the multi-grained output masks of SAM and iteratively introducing new prompts in regions where the student model exhibits inaccuracies. These tailored prompts guide the mask decoder, emphasizing areas of incorrect segmentation, thereby enhancing the learning process. Furthermore, we investigated various training configurations, examining different types of prompts, the impacts of freezing encoder/decoder components, the selection of distillation targets, etc. The specifics and findings of these explorations are detailed in the experimental section.
@@
In our final observations, we note that SAM, having been trained on a dataset with multi-grained annotations, encounters challenges in resolving the granularity of output when faced with ambiguous prompts, such as a single point. This is particularly evident when SAM is prompted with center points on the COCO dataset during evaluations; the model does not consistently produce instance-level masks, but rather part-level masks. This issue becomes more pronounced when SAM functions as the teacher model. To address this, we propose a simple yet effective module designed to explicitly discern and adapt to the granularity priors specific to a given test set or application scenario. This module enhances SAM’s ability to accurately interpret and respond to varying levels of prompt ambiguity.
Consequently, our EdgeSAM model achieves a remarkable performance boost, operating 37 times faster than the original SAM and 0.6 times more swiftly than MobileSAM and EfficientSAM on an NVIDIA 2080 Ti GPU. Notably, on the iPhone 14, EdgeSAM leverages its purely CNN-based backbone [50] to process image encoding in a mere 26 ms per image. This rate is over 7 times faster compared to the performance of MobileSAM and EfficientSAM on the same platform, marking it as the first SAM variant capable of real-time operation on edge devices. In terms of accuracy, EdgeSAM closely parallels the original SAM in box-prompt performance across the SA-1B [24], COCO [31], and LVIS [12] datasets and surpasses MobileSAM and EfficientSAM by an average of 1-2 mIoU across these datasets. This capability to run SAM in real-time on edge devices opens up a multitude of possibilities for downstream applications, including on-device video editing and video instance segmentation.
@@ 2 Related Work
Efficient Model Design. This research direction mainly focuses on designing efficient CNNs [18, 43, 17, 23, 65, 35, 13], transformers [36, 40, 37], and their mixed architecture [62, 34, 8, 25], with the goal of visual representation learning. Our work also adopts efficient models as the image encoder but is orthogonal to these works as it can be applied to various efficient backbones.
@@
Knowledge Distillation in Detection and Segmentation. The majority of research in knowledge distillation has been centered on classification tasks [14, 11, 56, 60, 66]. Several studies [55, 32, 44, 52, 63, 51] have applied knowledge distillation techniques to dense prediction tasks like semantic segmentation and object detection. Common approaches in these works involve leveraging pixel-wise correlations or channel-wise interactions between dense features of teacher and student models. Recently, there has been an increased interest in developing specialized knowledge distillation losses for query-based detectors such as DETR [4], as evidenced by works like [21, 7, 5]. MobileSAM [61], closely related to our work, implements pixel-wise feature distillation between the SAM encoder and a compact backbone. However, it does not address the prompt encoder and mask decoder, leading to a significant performance discrepancy compared to the original SAM. FastSAM [68], on the other hand, trains a YOLACT-based [2] instance segmentation model using the SA-1B dataset and employs heuristic rules for post-process object selection, a method that aligns marginally with the SAM principles. A concurrent work of ours, EfficientSAM [57], has achieved a great speed-performance trade-off through masked image pre-training, but it consumes a huge computational cost during training and as it uses the same image encoder as MobileSAM, it runs no faster than MobileSAM. Our work further explores the setting where both training and inference budgets are more limited.
@@
Efficient Segmentation Models. Prior studies in efficient segmentation [67, 26, 59, 20, 15, 49, 58, 27, 39, 38, 16] have predominantly concentrated on close-set segmentation within specific domains, with a significant portion of this research [26, 20, 39] specifically targeting driving scenarios. More recently, a few works [64, 49] have ventured into designing segmentation models suitable for on-device implementation, capable of running efficiently on mobile platforms. However, the realm of on-device interactive segmentation remains largely unexplored. MobileSAM [61] represents an initial attempt, but it still encounters challenges regarding computational efficiency and a noticeable decline in performance, indicating a substantial opportunity for further advancements in this direction.
@@ 3.1 Preliminaries: SAM
The SAM model consists of three components: image encoder, prompt encoder, and mask decoder. In particular, the image encoder takes up the most computational cost and parameters. It follows the backbone design of ViTDet [28], which essentially is a hierarchical vision transformer. The input and output sizes of the image encoder are kept fixed at $(3,1024,1024)$ and $(256,64,64)$.
SAM can handle four types of prompts, including point, box, mask, and text. Points and boxes are considered as the sparse prompts, which are encoded with random positional embeddings [48] summed with a type indicator token. Masks and free-form texts are embedded with a tiny CNN and the text encoder of CLIP [41], respectively. In this paper, we focus on the point and box prompts.
@@ 3.2 EdgeSAM
The goal of EdgeSAM is to transfer the capabilities of SAM into a much more compact model, which makes deployment on edge devices feasible. In particular, as shown in Fig. 1, EdgeSAM maintains the encoder-decoder architecture of SAM and aims to preserve the performance of zero-shot interactive segmentation with box and point prompts. We train EdgeSAM with only 1% data from the SA-1B dataset and evaluate its zero-shot transferability on the COCO [31] and LVIS [12] datasets.
@@ 3.2.1 Encoder-Only Knowledge Distillation
Inspired by MobileSAM [61], we adopt a pixel-wise feature distillation loss $L_{p}$ between the image encoder $T_{enc}$ of SAM and an efficient network $S_{enc}$ as follows:
@@
where $I$ denotes the input image. Since the downsampling stride and feature channels of the student model and SAM image encoder are not aligned, MobileSAM removes the downsampling operations in the last two stages of the student model and uses a projection layer to align the channel dimension. We also use the projection layer for channel alignment, but we keep the downsampling layers unchanged. Instead, we construct a tiny FPN [29] that upsamples the feature to the desired resolution and performs element-wise addition with features from previous stages.
We explore various efficient backbones following this paradigm, including ViT-based [53], CNN-based [50], and hybrid networks [3]. However, we find there is always a considerable performance gap. Training with a longer schedule or using distillation losses [32, 44] designed for dense prediction tasks does not show an obvious improvement. Therefore, we further propose to consider prompts during distillation to provide task-specific guidance.
@@ 3.2.2 Prompt-In-the-Loop Knowledge Distillation
where $\phi(\cdot)$ is binary thresholding, $f_{t}$ and $f_{s}$ denote the feature from the teacher and student image encoder, respectively. The teacher and student share the same set of $p$, $m$, and $\mathbf{c}$, which are kept frozen during training. We adopt a combination of the Dice loss [46] and BCE loss as the mask loss $L_{\text{mask}}$. Note that we allow the gradient backpropagate to the image encoder so that it is jointly learned.
While the distillation loss is frustratingly simple, the prompt selection for each training iteration needs to be carefully designed. We observe that fine-tuning the mask decoder poses a threat to the zero-shot generalization ability. That is, training with a particular combination of prompts sabotages its capability when inferencing with prompt combinations that are not used during training. For instance, training only with points leads to a significant performance drop when testing with box prompts. While freezing the mask decoder or using LoRA [19] on the decoder for regulation mitigates the problem, it also limits the performance upper-bound for prompts observed during training. Meanwhile, we find that even the original SAM outputs unsatisfactory mask predictions when prompted with ambiguous prompts, such as a single point. As shown in Fig. 2, on the COCO dataset, with the object center as the prompt, the mask mIoU of SAM is only 53.6. Thus, aligning the student’s outputs in such circumstances might not be optimal.
@@
Drawing inspiration from recent advancements in interactive segmentation methods [45, 10], our strategy involves iteratively sampling new prompts in the loop during the distillation phase. We start with an initial prompt. With equal probability, either a box or point prompt, as provided by the SA-1B dataset, is input into the decoders of both the teacher and student models. Subsequently, we identify areas where the mask predictions from the teacher and student diverge. As shown in Fig. 2, using the teacher’s output as the reference, we uniformly sample new prompts: a positive point in regions marked as false negatives, or a negative point in areas identified as false positives. These newly sampled points are then amalgamated with the existing prompts for the subsequent iteration of decoding. Please refer to Appendix 0.D for the algorithmic pseudocode. It is important to note that each prompt leads to four mask predictions at varying levels of granularity. In our analysis, the disagreement is calculated specifically between the teacher mask with the highest IoU score and its corresponding student mask.

### v3

@@ 1 Introduction
In this study, we explore the feasibility of deploying the Segment Anything Model (SAM) (Kirillov et al, 2023) directly on edge devices, such as smartphones, to enable real-time interactive segmentation and facilitate its integration in various downstream tasks. Addressing this challenge is non-trivial and non-achievable with the usual engineering efforts of compressing the model. First, SAM was not designed for on-device deployment. It has a total parameter count of 641M and requires 2735 GFLOPs under its fixed $1024\times 1024$ input resolution. We failed to run SAM on an iPhone 14 due to its large computation and memory consumption. Even on an NVIDIA 2080 Ti, the throughput is just four images per second.
@@
A seemingly straightforward solution to speed up is to replace SAM’s huge ViT-based image encoder with a more compact version, as demonstrated by MobileSAM (Zhang et al, 2023a). This approach indeed leads to a considerable speed boost, approximately 23 times faster, but it also substantially reduces the performance. For instance, the mask mIoU on the COCO dataset (Lin et al, 2014) drops from 77.3 to 74.4 with ground-truth boxes as the box prompts. Moreover, when deployed on edge devices, the speed of MobileSAM is still far from being real-time. For example, on iPhone 14, its throughput is only five images per second. A more recent work, EfficientSAM (Xiong et al, 2023) follows the same architecture as MobileSAM but further leverages the masked image pre-training to boost the performance. However, it runs at the same speed as MobileSAM and consumes huge computational costs during training.
@@
In our extensive empirical investigations, it became unexpectedly clear that the key factor in knowledge distillation from SAM is not the incorporation of losses specifically tailored for dense tasks (Liu et al, 2019; Shu et al, 2021) or query-based detectors (Huang et al, 2022; Chen et al, 2022a; Chang et al, 2023). Rather, it is the strategic selection of prompts during the distillation process that holds paramount importance. Addressing this, we introduce a novel approach: a dynamic prompt-in-the-loop strategy. This technique effectively enables the student model to assimilate the intricate knowledge encapsulated within SAM. It involves actively aligning the student model with the multi-grained output masks of SAM and iteratively introducing new prompts in regions where the student model exhibits inaccuracies. These tailored prompts guide the mask decoder, emphasizing areas of incorrect segmentation, thereby enhancing the learning process. Furthermore, we investigated various training configurations, examining different types of prompts, the impacts of freezing encoder/decoder components, the selection of distillation targets, etc.. The specifics and findings of these explorations are detailed in the experimental section.
@@
In our final observations, we note that SAM, having been trained on a dataset with multi-grained annotations, encounters challenges in resolving the granularity of output when faced with ambiguous prompts, such as a single point. This is particularly evident when SAM is prompted with center points on the COCO dataset during evaluations; the model does not consistently produce instance-level masks, but rather part-level masks. This issue becomes more pronounced when SAM functions as the teacher model. To address this, we propose a simple yet effective module designed to explicitly discern and adapt to the granularity priors specific to a given test set or application scenario. This module enhances SAM’s ability to interpret and respond to varying levels of prompt ambiguity accurately.
Consequently, our EdgeSAM model achieves a remarkable performance boost, operating 37 times faster than the original SAM and 0.6 times more swiftly than MobileSAM and EfficientSAM on an NVIDIA 2080 Ti GPU. Notably, on the iPhone 14, EdgeSAM leverages its purely CNN-based backbone (Wang et al, 2023b) to process image encoding in a mere 26 ms per image. This rate is over 7 times faster compared to the performance of MobileSAM and EfficientSAM on the same platform, marking it as the first SAM variant capable of real-time operation on edge devices. In terms of accuracy, EdgeSAM closely parallels the original SAM in box-prompt performance across the SA-1B (Kirillov et al, 2023), COCO (Lin et al, 2014), and LVIS (Gupta et al, 2019) datasets and surpasses MobileSAM and EfficientSAM by an average of 1-2 mIoU across these datasets. This capability to run SAM in real-time on edge devices opens up a multitude of possibilities for downstream applications, including on-device video editing and video instance segmentation.
@@ 2 Related Work
Efficient Model Design. This research direction mainly focuses on designing efficient CNNs (Howard et al, 2017; Sandler et al, 2018; Howard et al, 2019; Iandola et al, 2016; Zhang et al, 2018; Ma et al, 2018; Han et al, 2020), transformers (Maaz et al, 2022; Pan et al, 2022; Mehta and Rastegari, 2022), and their mixed architecture (Zhang et al, 2023b; Ma et al, 2022; Chen et al, 2022b; Li et al, 2022a), with the goal of visual representation learning. Our work also adopts efficient models as the image encoder but is orthogonal to these works as it can be applied to various efficient backbones.
Figure 1: Figure (a) shows the overview of EdgeSAM. We first apply encoder-only knowledge distillation between the output feature of the image encoder of SAM and EdgeSAM. Then, we adopt prompt-in-the-loop knowledge distillation, which interactively samples point prompts from the wrongly segmented areas. Note that the initial prompt can also be in a box. In (b), we show its throughput and the mIoU performance on the COCO and LVIS datasets with box prompts compared with SAM, MobileSAM, and EfficientSAM.
@@
Knowledge Distillation in Detection and Segmentation. The majority of research in knowledge distillation has been centered on classification tasks (Hinton et al, 2014; Guan et al, 2020; Xie et al, 2020; Yuan et al, 2020; Zhang et al, 2020). Several studies (Xie et al, 2018; Liu et al, 2019; Shu et al, 2021; Wang et al, 2020; Zhang and Ma, 2020; Wang et al, 2019) have applied knowledge distillation techniques to dense prediction tasks like semantic segmentation and object detection. Common approaches in these works involve leveraging pixel-wise correlations or channel-wise interactions between dense features of teacher and student models. Recently, there has been an increased interest in developing specialized knowledge distillation losses for query-based detectors such as DETR (Carion et al, 2020), as evidenced by works like (Huang et al, 2022; Chen et al, 2022a; Chang et al, 2023). MobileSAM (Zhang et al, 2023a), closely related to our work, implements pixel-wise feature distillation between the SAM encoder and a compact backbone and works following this direction include EfficientViT-SAM (Zhang et al, 2024), RepViT-SAM (Wang et al, 2023a), and NanoSAM (nan, 2023). However, they do not address the prompt encoder and mask decoder, leading to a significant performance discrepancy compared to the original SAM. In particular, RepViT-SAM follows MobileSAM and replaces its image encoder with RepViT, thus yielding the same speed as EdgeSAM. However, because it only adopts the encoder-only KD, its performance falls behind our work. FastSAM (Zhao et al, 2023), on the other hand, trains a YOLACT-based (Bolya et al, 2019) instance segmentation model using the SA-1B dataset and employs heuristic rules for post-process object selection, a method that aligns marginally with the SAM principles. EfficientSAM (Xiong et al, 2023) has achieved a great speed-performance trade-off through masked image pre-training, but it consumes a huge computational cost during training, and as it uses the same image encoder as MobileSAM, it runs no faster than MobileSAM. Our work further explores the setting where both training and inference budgets are more limited.
Recently, SAM 2 (Ravi et al, 2024) extends SAM to the video domain, and several works optimize SAM 2 for efficient inference (Xiong et al, 2024; Zhou et al, 2025). Our work can also contribute to the SAM pre-training stage for these methods. The design philosophy of the proposed prompt-in-the-loop knowledge distillation is also related to refinement-based dense prediction methods such as Cascade-RCNN Cai and Vasconcelos (2018) and RefineMask Zhang et al (2021), but with a focus on distillation instead of architecture.
@@
Efficient Segmentation Models. Prior studies in efficient segmentation (Zhao et al, 2018; Li et al, 2020; Yu et al, 2018; Hu et al, 2023; Hong et al, 2021a; Wan et al, 2023; Yu et al, 2021; Li et al, 2023; Mehta et al, 2019, 2018; Hong et al, 2021b) have predominantly concentrated on close-set segmentation within specific domains, with a significant portion of this research (Li et al, 2020; Hu et al, 2023; Mehta et al, 2019) specifically targeting driving scenarios. More recently, a few works (Zhang et al, 2022; Wan et al, 2023) have ventured into designing segmentation models suitable for on-device implementation, capable of running efficiently on mobile platforms. However, the realm of on-device interactive segmentation remains largely unexplored. MobileSAM (Zhang et al, 2023a) represents an initial attempt, but it still encounters challenges regarding computational efficiency and a noticeable decline in performance, indicating a substantial opportunity for further advancements in this direction. More related research can be found in this survey (Sun et al, 2024).
@@ 3.1 Preliminaries: SAM
The SAM model consists of three components: image encoder, prompt encoder, and mask decoder. In particular, the image encoder takes up the most computational cost and parameters. It follows the backbone design of ViTDet (Li et al, 2022b), which essentially is a hierarchical vision transformer. The input and output sizes of the image encoder are kept fixed at $(3,1024,1024)$ and $(256,64,64)$.
SAM can handle four types of prompts, including point, box, mask, and text. Points and boxes are considered as the sparse prompts, which are encoded with random positional embeddings (Tancik et al, 2020) summed with a type indicator token. Masks and free-form texts are embedded with a tiny CNN and the text encoder of CLIP (Radford et al, 2021), respectively. In this paper, we focus on the point and box prompts. Note that, when no prompt is given to SAM, SAM can operate in the segment everything mode, where it generates a grid of positive points and regards each point as a prompt.
@@ 3.2 EdgeSAM
The goal of EdgeSAM is to transfer the capabilities of SAM into a much more compact model, which makes deployment on edge devices feasible. In particular, as shown in Fig. 1, EdgeSAM maintains the encoder-decoder architecture of SAM and aims to preserve the performance of zero-shot interactive segmentation with box and point prompts. We train EdgeSAM with only 1% data from the SA-1B dataset and evaluate its zero-shot transferability on the COCO (Lin et al, 2014) and LVIS (Gupta et al, 2019) datasets.
@@ 3.2.1 Encoder-Only Knowledge Distillation
Inspired by MobileSAM (Zhang et al, 2023a), we adopt a pixel-wise feature distillation loss $L_{p}$ between the image encoder $T_{enc}$ of SAM and an efficient network $S_{enc}$ as follows:
@@
where $I$ denotes the input image. Since the downsampling stride and feature channels of the student model and SAM image encoder are not aligned, MobileSAM removes the downsampling operations in the last two stages of the student model and uses a projection layer to align the channel dimension. We also use the projection layer for channel alignment, but we keep the downsampling layers unchanged. Instead, we construct a tiny FPN (Lin et al, 2017a) that upsamples the feature to the desired resolution and performs element-wise addition with features from previous stages. Specifically, the FPN takes the output feature maps from the last two stages of the backbone as the input and projects them to the same number of channels as SAM, followed by upsampling and addition as standard FPN operations.
We explore various efficient backbones following this paradigm, including ViT-based (Wu et al, 2022), CNN-based (Wang et al, 2023b), and hybrid networks (Cai et al, 2023). However, we find there is always a considerable performance gap. Training with a longer schedule or using distillation losses (Liu et al, 2019; Shu et al, 2021) designed for dense prediction tasks does not show an obvious improvement. Therefore, we further propose to consider prompts during distillation to provide task-specific guidance.
@@ 3.2.2 Prompt-In-the-Loop Knowledge Distillation
where $\phi(\cdot)$ is binary thresholding, $f_{t}$ and $f_{s}$ denote the feature from the teacher and student image encoder, respectively. The teacher and student share the same set of $p$, $m$, and $\mathbf{c}$, which are kept frozen during training. We adopt a combination of the Dice loss (Sudre et al, 2017) and BCE loss as the mask loss $L_{\text{mask}}$. Note that we allow the gradient backpropagate to the image encoder so that it is jointly learned.
While the distillation loss is frustratingly simple, the prompt selection for each training iteration needs to be carefully designed. We observe that fine-tuning the mask decoder poses a threat to the zero-shot generalization ability. That is, training with a particular combination of prompts sabotages its capability when inferencing with prompt combinations that are not used during training. For instance, training only with points leads to a significant performance drop when testing with box prompts. While freezing the mask decoder or using LoRA (Hu et al, 2022) on the decoder for regulation mitigates the problem, it also limits the performance upper-bound for prompts observed during training. Meanwhile, we find that even the original SAM outputs unsatisfactory mask predictions when prompted with ambiguous prompts, such as a single point. As shown in Fig. 2, on the COCO dataset, with the object center as the prompt, the mask mIoU of SAM is only 53.6. Thus, aligning the student’s outputs in such circumstances might not be optimal.
$T_{enc},S_{enc}\leftarrow\text{SAM / EdgeSAM encoder}$;
$T_{dec}\leftarrow\text{SAM decoder}$;
$S_{dec}\leftarrow\text{EdgeSAM decoder initialized w/}\ T_{dec}$;
$m,\mathbf{c}\leftarrow\text{shared mask / IoU tokens}$;
$\mathcal{I},\mathcal{P}\leftarrow\text{images and prompts for training}$;
$N,M\leftarrow\text{training steps, prompt sampling loops}$;
for $i=1,2,\dots,N$ do
 $f_{t},f_{s}\leftarrow T_{enc}(\mathcal{I}_{i}),S_{enc}(\mathcal{I}_{i})$;
 $p\leftarrow\text{select the box or point prompt in}\ \mathcal{P}_{i}$;
 $L\leftarrow L_{\text{mask}}(m_{t},m_{s})$;
 for $j=1,2,\dots,M$ do
    $\hat{p}\leftarrow\text{sample\_in\_disagree}(m_{t},m_{s})$;
    pp^ $\leftarrow\hat{p}\text{ appends to}\ p$;
    $L\leftarrow L+L_{\text{mask}}(m_{t},m_{s})$;
   end for
 $S_{enc},S_{dec}\leftarrow\text{SGD model update}$;
end for
Algorithm 1 Prompt-In-the-Loop Distillation
@@
Drawing inspiration from recent advancements in interactive segmentation methods (Sofiiuk et al, 2022; Forte et al, 2020), our strategy involves iteratively sampling new prompts in the loop during the distillation phase. As detailed in Algo. 1, we start with an initial prompt. With equal probability, either a box or point prompt, as provided by the SA-1B dataset, is input into the decoders of both the teacher and student models. Subsequently, we identify areas where the mask predictions from the teacher and student diverge. As shown in Fig. 2, using the teacher’s output as the reference, we uniformly sample new prompts: a positive point in regions marked as false negatives, or a negative point in areas identified as false positives. These newly sampled points are then amalgamated with the existing prompts for the subsequent iteration of decoding. Note that we cannot sample new box prompts from the disagreement area because the original SAM only supports up to one box for each object of interest. It is important to note that each prompt leads to four mask predictions at varying levels of granularity. In our analysis, the disagreement is calculated specifically between the teacher mask with the highest IoU score and its corresponding student mask.